OpenStudy (anonymous):

The function g is defined by g(x) = sin x, for 0≤x<1/2π px-1, for 1/2π≤x<2π q , for x≥2π where p and q are constants. Find the values of p and q if g is continuous at x=1/2π and x=2π

hartnn (hartnn):

if g is continuous at x=1/2π then \(\Large \lim \limits_{x\rightarrow \pi/2^-}g(x) =\lim \limits_{x\rightarrow \pi/2^+}g(x)\)

hartnn (hartnn):

what will u choose for \(\pi/2^-\) ? sin x or px-1 ?

OpenStudy (anonymous):

px-1 maybe

hartnn (hartnn):

let me tell you what \(\pi/2^-\) it is a number which is very near to pi/2 but LESS then pi/2 so, you will choose the function which has \(x<\pi/2\)

hartnn (hartnn):

now once again, what will u choose for π/2− ? sin x or px-1 ?

OpenStudy (anonymous):

sin x :D

hartnn (hartnn):

correct! :) so you will choose px-1 for x->pi/2+ now can u find those 2 limits ? in my first comment ?

OpenStudy (anonymous):

can you write the all solution please ..

hartnn (hartnn):

\(\Large \lim \limits_{x\rightarrow \pi/2^-}g(x) =\lim \limits_{x\rightarrow \pi/2^+}g(x) \\ \Large \lim \limits_{x\rightarrow \pi/2^-}\sin x =\lim \limits_{x\rightarrow \pi/2^+}px-1\) you know how to solve limits questions, right ?

OpenStudy (anonymous):

wait i try it first

hartnn (hartnn):

sure, go ahead :) take your time....

OpenStudy (anonymous):

i got the wrong answer .. the correct answer is p=4/π but i got 2.055/π .. what should i do ..

hartnn (hartnn):

did u put x = pi/2 in sin x and px-1 ?

hartnn (hartnn):

what is sin pi/2 = .... ?

OpenStudy (anonymous):

0.027

hartnn (hartnn):

no! you used your calculator in degree mode! you need to use it in radians mode :) because pi/2 is in radians

hartnn (hartnn):

sin pi/2 = 1 actually.

hartnn (hartnn):

so, now can you get p from 1 = p (pi/2) -1 ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok i got the correct answer for p ,, i use wrong mode in calculator .. can u give me the other equations for find q ?

hartnn (hartnn):

sure,

hartnn (hartnn):

\(\Large \lim \limits_{x\rightarrow 2\pi^-}g(x) =\lim \limits_{x\rightarrow 2\pi^+}g(x) \\ \Large \lim \limits_{x\rightarrow 2\pi^-}px-1 =\lim \limits_{x\rightarrow 2\pi^+}q\)

OpenStudy (anonymous):

ok i try it now ..

hartnn (hartnn):

when you are done, tell me what u got for q

OpenStudy (anonymous):

i got q= 7

hartnn (hartnn):

good work! i got 7 too :)

OpenStudy (anonymous):

Thank a lot sir .. :D

hartnn (hartnn):

welcome ^_^ and since you are new here \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{ }}\color{#e6e600}{\text{S}}\color{green}{\text{t}}\color{blue}{\text{u}}\color{purple}{\text{d}}\color{purple}{\text{y}}\color{red}{\text{!}}\color{orange}{\text{!}}\color{#e6e600}{\text{ }}\color{green}{\text{:}}\color{blue}{\text{)}}\color{purple}{\text{}}\end{array} \]

OpenStudy (anonymous):

Thank For helping me .. if i got any problem's and i 'm sure i will discuss in OPEN STUDY .. Such A Good website for student like me ^_^

hartnn (hartnn):

i am glad to hear that! :) its good for everyone ;)